OpenStudy (anonymous):

Determine whether the graph of y = x^2 + 4x - 7 has a maximum or minimum point, then find the maximum or minimum value. Answer Minimum; (-11, -2) Maximum; (-11, -2) Maximum; (-2, -11) Minimum; (-2, -11)

OpenStudy (anonymous):

hey can you help me i got like 15 questions i need help on then I know the rest on my paper

OpenStudy (cwrw238):

x^2 + 4x - 7 = (x + 2)^2 - 4 - 7 = (x+2)^2 - 11 when x = -2 what is value of (x+2)^2 - 11?

OpenStudy (cwrw238):

equation with positive x^2 term have a minimum - now you should be able to pick the right option

OpenStudy (cwrw238):

sorry - gotta go now